### v3

@@ 3.2 Avoidance of strict saddle point
The map $S$ has the property of so-called “active manifold identification”, which means the iterates fall onto a smooth manifold after finite iterations. The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [33, 32] and the manifold in this case is simply the subspace consists of the nonzeros at a limit point.
@@
    $S$ is $C^{1}$-smooth in $B$ and $S^{x}(x,\epsilon)=x-\frac{1}{\beta}g(x,\epsilon)$ with
@@
(i) Since $S$ is a continuous map and $x^{*}=S^{x}(x^{*},\epsilon^{*})$, it holds that
@@
(ii) It follows from (i) that there exists a neighborhood $B$ around $(x^{*},\epsilon^{*})$, so that the subproblem at $(x^{k},\epsilon^{k})\in B$ agree with following constrain problem
@@
Therefore, $S$ is $C^{1}$-smooth in $B$
(iii) It follows from (27) that the iteration $T$ is $C^{1}$-smooth in $B$ for any for any $(x,\epsilon)\in B$. ∎
The above theorem shows the local behavior of ${\rm DIRL}_{1}$: it is equivalent to the gradient descent for the variables in $\mathcal{I}^{*}$ locally around a stationary point in $x^{*}+\mathcal{M}_{\mathcal{I}^{*}}$. Specifically, it can be expressed as
@@
around $(x^{*},\epsilon^{*})$. We then proceed to the saddle escape mechanism for DIRL1, which ensures that the algorithm does not converge to the strict saddle points defined in Definition 3 and, instead, converges to local minima. To apply the CSM theorem, we have to show the unstable fixed point of $T$ corresponds to the strict saddle point of (P).
@@
To apply Corollary 1, we next to show $T$ is a lipeomorphism, even if it might not be differentiable globally on $\mathbb{R}^{2n}$. This is equivalent to show $T$ and its inverse $T^{-1}$ is Lipschitz continuous.
@@
(Lipeomorphism) Suppose Assumption 1-3 hold. Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C))$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ defined in Proposition 10, and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$. Then $T$ is invertible, and $T$ and $T^{-1}$ is Lipschitz continuous (i.e., $T$ is a lipeomorphism) when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
@@
If $r^{\prime}(|0+|)=\infty$, by Proposition 10, $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C$, $\forall x\in\mathcal{L}(F^{0})$. When $r^{\prime}(|x_{i}|)\geq w(x_{i},\epsilon_{i})>C$, then it must be true that $S_{i}(x,\epsilon)=0$. Therefore, for any $S_{i}(x)=0$ with $x\in\mathcal{L}(F^{0})$,
@@
by Assumption 3.
This implies that we can equivalently write $S$ as
$S_{i}(x,\epsilon)=\mathbb{S}_{\lambda u_{i}/\beta}(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$
@@
Next we prove $S$ is Lipschitz continuous. It follows that
@@
where the first inequality is by Lemma A.3 and the second is by (19) and (30). Therefore, $\|\alpha(\hat{x}-S^{x}(\hat{x},\hat{\epsilon}))-\alpha(\tilde{x}-S^{x}(\tilde{x},\tilde{\epsilon}))\|\leq\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))$ holds for any $\hat{y},\tilde{y}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$.
@@
(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumption 1-3. Let the condition of Proposition 14 holds and $\alpha<\beta/\rho$. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10. Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.
@@
It follows directly from Theorem 13 that map $T$ is local $C^{1}$-smooth around any stationary points, and all stationary points $x^{*}$ are unstable fixed points. By applying Proposition 14, $T$ is a lipeomorphism. By Corollary 1, the set of initial conditions attracted by such fixed points has zero Lebesgue measure. ∎
We illustrate this property by applying DIRL1 to Example 1. We test $\alpha=0.1$ and $0.9$ with different initial points; first is $(3,3)$ representing an initial point far from the global minimizer and the second is $(0,0.043)$ very close to the saddle point. The trajectories are shown in Figure 3. In four cases, the algorithm always converges to the global minimizer $(0,1)$.
@@ 4 Damped Iterative Reweighted $\ell_{2}$ Algorithms
The ${\rm IRL}_{2}$ algorithm is an effective method to solve problem (P). In this method, we write $h(\cdot)=r\circ\sqrt{(\cdot)}$, so that $s(x_{i})=h(x^{2}_{i})$. At each iteration, after adding perturbation $\epsilon_{i}^{2}$ to each $h$ and linearize $h$ at $(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}$, the ${\rm IRL}_{2}$ algorithm solves the following weighted $\ell_{2}$ regularized subproblem,
@@
The subproblem solution is given by $y^{k}=S^{x}(x^{k},\epsilon^{k})$, where the operator $S$ is defined as
$S_{i}^{x}(x,\epsilon):=\frac{1}{1+\frac{2\lambda}{\beta}w(x_{i},\epsilon_{i})}(x_{i}-\frac{1}{\beta}\nabla_{i}f(x)).$ (31)
If $w(x_{i},\epsilon_{i})=\infty$, we simply set $S_{i}^{x}(x,\epsilon)=0$. The damped iteratively reweighted $\ell_{2}$ algorithm, hereinafter named DIRL2 algorithm, is presented in Algorithm 10.
@@
The fixed point iteration of DIRL2 is then given as $(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=(1-\alpha)I+\alpha S$, where
@@
Suppose $x^{*}$ is a stationary point. It follows from (6) that for $i\in\mathcal{I}^{*}$, $x_{i}^{*}=x_{i}^{*}-\tfrac{1}{\beta}(\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*}))$ implying $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$. Therefore, $x_{i}^{*}=S_{i}(x^{*},0)$ since $w(x_{i}^{*},0)=\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}$. As for $i\in\mathcal{J}^{*}$, $w(0,0)=\infty$, implying that $0=S_{i}(x^{*},0)$. To summarize, $x^{*}=S^{x}(x^{*},0)$.
@@
The converse statement of Proposition 16 may not be true when $\mathcal{J}^{*}\neq\emptyset$ and $r^{\prime}(|0^{+}|)<\infty$. Since for any $x^{*}$ satisfying $x_{i}^{*}=S_{i}(x^{*},0)$, $i\in\mathcal{I}^{*}$, it holds naturally true that $w(0,0)=\infty$ and $S_{i}(x^{*},0)=0$. However, there is no guarantee that $x^{*}$ satisfies (7). In those cases, $S$ may have two types of fixed points—those are stationary points and others are not. This represents a stark difference from the DIRL1 algorithm, where all the fixed points of the iteration map are stationary points.
@@ 4.1 Convergence analysis
In this subsection, we show that DIRL2 only converge to its fixed points which are the stationary points of (P). Notice that this is true in all cases even though it may have fixed points that are not the stationary points, as discussed in the previous subsection.
Similar to the analysis of DIRL1 algorithm, the analysis of DIRL2 is also based on the monotonic decrease of the objective $F(x,\epsilon)$. We also need Assumption 3 and consequently (19) hold true for the analysis. However, it must be noticed that the definition of $F(x,\epsilon)$ is different from that for DIRL1 algorithm.
@@
Since $h(\cdot)=r\circ\sqrt{(\cdot)}$ is also concave by Assumption 1, it follows that for $i=1,...,n$,
@@
where the first inequality is by Lemma A.2 and the second is by the convexity of $(\cdot)^{2}$. Combine (24), (35) and get
@@
which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, it follows from the boundedness of $\{x^{k}\}$ that there must exist $C>0$ such that (21) holds for any $k\in\mathbb{N}$. ∎
@@
(Global convergence) Suppose Assumption 1 and 3 hold. Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{2}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$. Then $\mu^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is first-order optimal for (P).
@@
Firstly, boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. The optimality condition of the $k$th subproblem implies
$-\beta(y_{i}^{k}-x_{i}^{k})=\nabla_{i}f(x^{k})+2\lambda w(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}=\nabla_{i}f(x^{k})+\frac{r^{\prime}(\sqrt{x_{i}^{2}+\epsilon_{i}^{2}})}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}y_{i}^{k}.$ (36)
As $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$, $y_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$, $\lim\limits_{k\in\mathcal{S}}\frac{y_{i}^{k}}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}=\text{sign}(x_{i}^{*})$ for $i\in\mathcal{I}^{*}$, and $\limsup\limits_{k\in\mathcal{S}}\frac{|y_{i}^{k}|}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}\leq 1$ for $i\in\mathcal{J}^{*}$.
Therefore, taking $k\to\infty,k\in\mathcal{S}$ in both sides of (36), we have for $i\in\mathcal{I}^{*}$
$\nabla_{i}f(x^{*})+2\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*})=0,$
satisfying (6). As for $i\in\mathcal{J}^{*}$ and $r^{\prime}(|0^{+}|)<\infty$,
$|\nabla_{i}f(x^{*})|=\lim_{k\in\mathcal{S}\atop k\to\infty}|\nabla_{i}f(x^{k})|\leq\lim_{k\in\mathcal{S}\atop k\to\infty}|\frac{r^{\prime}(\sqrt{x_{i}^{2}+\epsilon_{i}^{2}})}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}y_{i}^{k}|=\limsup\limits_{k\in\mathcal{S}\atop k\to\infty}\frac{r^{\prime}(|0^{+}|)|y_{i}^{k}|}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}\leq r^{\prime}(|0^{+}|),$
satisfying (7) by Assumption 2. Overall, $x^{*}$ is stationary. ∎

### v4

@@ 3.2 Avoidance of strict saddle point
The map $S$ has the property of so-called “active manifold identification”, which means the iterates fall onto a smooth manifold after finite iterations. The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [33, 32] and the manifold in this case is simply the subspace consisting of nonzeros at a limit point.
@@
    $S$ is $C^{1}$ smooth in $B$ and $S^{x}(x,\epsilon)=x-\frac{1}{\beta}g(x,\epsilon)$ with
@@
(i) Since $S$ is a continuous map and $x^{*}=S^{x}(x^{*},\epsilon^{*})$, we have
@@
(ii) It follows from (i) that there exists a neighborhood $B$ around $(x^{*},\epsilon^{*})$, so that the subproblem at $(x^{k},\epsilon^{k})\in B$ agrees with following constrained problem
@@
Therefore, $S$ is $C^{1}$ smooth in $B$.
(iii) It follows from (29) that the iteration $T$ is $C^{1}$ smooth in $B$. ∎
The above theorem shows the local behavior of ${\rm DIRL}_{1}$: it is equivalent to the gradient descent method for the variables in $\mathcal{I}^{*}$ locally around a stationary point in $x^{*}+\mathcal{M}_{\mathcal{I}^{*}}$. Specifically, it can be expressed as
@@
around $(x^{*},\epsilon^{*})$. We then proceed to the saddle escape mechanism for DIRL1, which ensures that the algorithm does not converge to the strict saddle points defined in Definition 3 and, instead, converges to local minima. To apply the CSM theorem, we have to show that the unstable fixed point of $T$ corresponds to the strict saddle point of (P).
@@
To apply Corollary 1, we next show that $T$ is a lipeomorphism, even if it might not be differentiable globally on $\mathbb{R}^{2n}$. This is equivalent to show $T$ and its inverse $T^{-1}$ is Lipschitz continuous.
@@
(Lipeomorphism) Suppose Assumption 1-3 hold. Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C))$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ is defined in Proposition 10, and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$. Then $T$ is invertible; moreover, $T$ and $T^{-1}$ is Lipschitz continuous (that is, $T$ is a lipeomorphism) when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
@@
If $r^{\prime}(|0+|)=\infty$, by Proposition 10, $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C$, $\forall x\in\mathcal{L}(F^{0})$. When $r^{\prime}(|x_{i}|)\geq w(x_{i},\epsilon_{i})>C$, it must be true that $S_{i}(x,\epsilon)=0$. Therefore, for any $S_{i}(x)=0$ with $x\in\mathcal{L}(F^{0})$,
@@
by Assumption 3. This implies that we can equivalently write $S$ as
$S_{i}(x,\epsilon)=\mathbb{S}_{\lambda u_{i}(x_{i},\epsilon_{i})/\beta}(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$
@@
Next we prove that $S$ is Lipschitz continuous. It follows that
@@
where the first inequality is determined by Lemma A.3 and the second is by (21) and (32). Therefore, $\|\alpha(\hat{x}-S^{x}(\hat{x},\hat{\epsilon}))-\alpha(\tilde{x}-S^{x}(\tilde{x},\tilde{\epsilon}))\|\leq\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))$ holds for any $\hat{y},\tilde{y}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$.
@@
(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumption 1-3 satisfied. Suppose $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$, $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$ and $\alpha<\beta/\rho$. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10. Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.
@@
It follows directly from Theorem 13 that the map $T$ is local $C^{1}$ smooth around any stationary points, and all stationary points $x^{*}$ are unstable fixed points. By applying Proposition 14, $T$ is a lipeomorphism. By Corollary 1, the set of initial conditions attracted by such fixed points has a zero Lebesgue measure. ∎
We illustrate this property by applying DIRL1 to Example 1. We test $\alpha=0.1$ and $0.9$ with different initial points: $(3,3)$ representing an initial point far from the global minimizer and $(0,0.043)$ very close to the saddle point. The trajectories are shown in Figure 3. In four cases, the algorithm always converges to the global minimizer $(0,1)$.
@@ 4 Damped Iterative Reweighted $\ell_{2}$ Algorithms
The ${\rm IRL}_{2}$ algorithm is an effective method to solve problem (P). In this method, we write $h(\cdot)=r\circ\sqrt{(\cdot)}$, so that $s(x_{i})=h(x^{2}_{i})$. Since $\sqrt{\cdot}$ is concave on $\mathbb{R}_{+}$, $\sqrt{\lambda y_{1}+(1-\lambda)y_{2}}\geq\lambda\sqrt{y_{1}}+(1-\lambda)\sqrt{y_{2}}$ for any $\lambda\in[0,1]$ and $y_{1},y_{2}\geq 0$. Since $r$ is increasing and concave, it follows that $h(\lambda y_{1}+(1-\lambda)y_{2})=r(\sqrt{\lambda y_{1}+(1-\lambda)y_{2}})\geq r(\lambda\sqrt{y_{1}}+(1-\lambda)\sqrt{y_{2}})\geq\lambda r(\sqrt{y_{1}})+(1-\lambda)r(\sqrt{y_{2}})=\lambda h(y_{1})+(1-\lambda)h(y_{2})$, meaning $h$ is also concave on $\mathbb{R}_{+}$.
At each iteration, after adding perturbation $\epsilon_{i}^{2}$ to each $h$ and linearize $h$ at $(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}$, the ${\rm IRL}_{2}$ algorithm solves the following weighted $\ell_{2}$ regularized subproblem (for brievty, we still use $S$ and $T$ for the maps),
@@
The solution of the subproblem is given by
$y^{k}_{i}=S_{i}^{x}(x,\epsilon):=\frac{1}{1+\frac{2\lambda}{\beta}w(x_{i}^{k},\epsilon_{i}^{k})}(x_{i}^{k}-\frac{1}{\beta}\nabla_{i}f(x^{k})).$ (33)
If $w(x_{i}^{k},\epsilon_{i}^{k})=\infty$, we simply set $S_{i}^{x}(x^{k},\epsilon^{k})=0$. The damped iteratively reweighted $\ell_{2}$ algorithm, hereinafter named DIRL2 algorithm, is presented in Algorithm 10.
@@
The fixed point iteration of DIRL2 is then given as $(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})$ with $T=(1-\alpha)I+\alpha S$ and
@@
Suppose $x^{*}$ is a stationary point. It follows from (6) that for $i\in\mathcal{I}^{*}$, $x_{i}^{*}=x_{i}^{*}-\tfrac{1}{\beta}(\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*}))$ implying $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$. Therefore, $x_{i}^{*}=S_{i}(x^{*},0)$ since $w(x_{i}^{*},0)=\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}$. As for $i\in\mathcal{J}^{*}$, $w(0,0)=\infty$, implying that $0=S_{i}(x^{*},0)$. In summary, $x^{*}=S^{x}(x^{*},0)$.
@@
The converse statement of Proposition 16 may not be true when $\mathcal{J}^{*}\neq\emptyset$ and $r^{\prime}(|0^{+}|)<\infty$. Since for any $x^{*}$ that satisfies $x_{i}^{*}=S_{i}(x^{*},0)$, $i\in\mathcal{I}^{*}$, it is naturally true that $w(0,0)=\infty$ and $S_{i}(x^{*},0)=0$. However, there is no guarantee that $x^{*}$ satisfies (7). In those cases, $S$ may have two types of fixed points — those are stationary points and others are not. This represents a stark difference from the DIRL1 algorithm, where all fixed points of the iteration map are stationary points.
@@ 4.1 Convergence analysis
In this subsection, we show that DIRL2 only converges to its fixed points which are the stationary points of (P) under the assumption that $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$. Similarly to the analysis of the DIRL1 algorithm, the analysis of DIRL2 is also based on the monotonic decrease in the objective $F(x,\epsilon)$. We also need Assumption 3 and consequently (21) to hold for the analysis. However, it should be noted that the definition of $F(x,\epsilon)$ is different from that of the DIRL1 algorithm.
@@
It follows from the concavity of $h$ that for $i=1,...,n$,
@@
where the first inequality is by Lemma A.2 and the second is by the convexity of $(\cdot)^{2}$. Combining (26) and (37), we get
@@
which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, from the boundedness of $\{x^{k}\}$, there must exist $C>0$ such that (23) holds for any $k\in\mathbb{N}$. ∎
@@
(Global convergence) Suppose Assumption 13 hold, and $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$. Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{2}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$. Then any $x^{*}\in\Omega$ is first-order optimal for (P).
@@
Firstly, the boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. The optimality condition of the $k$th subproblem implies
$-\beta(y_{i}^{k}-x_{i}^{k})=\nabla_{i}f(x^{k})+2\lambda w(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}=\nabla_{i}f(x^{k})+\frac{r^{\prime}(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})}{\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}}}y_{i}^{k}.$ (38)
As $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$, $y_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$, $\lim\limits_{k\in\mathcal{S}}\frac{y_{i}^{k}}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}=\text{sign}(x_{i}^{*})$ for $i\in\mathcal{I}^{*}$.
For $i\in\mathcal{J}^{*}$, there is nothing to prove since $|\nabla_{i}f(x^{*})|\leq r^{\prime}(|0^{+}|)$ naturally holds. In general, $x^{*}$ is stationary. ∎